OpenStudy (anonymous):

solve integral (1-tan^2(x))

OpenStudy (anonymous):

\[\int\limits_{?}^{?}1-\tan ^{2}x=\int\limits_{?}^{?}1-(\sec ^{2}x-1)\]

OpenStudy (yuki):

1-tan^2(x) is a nice Identity... do you know what it is ?

OpenStudy (anonymous):

\[=\int\limits_{?}^{?}2-\sec ^{2}x\]

OpenStudy (anonymous):

2x-tanx+C

OpenStudy (yuki):

just to let you know uzma, if you eliminate the _{}-{} part after "int" the limits of the integral will go away like this \[\int\limits\]

OpenStudy (anonymous):

oh thanks, i dint try it :)

OpenStudy (yuki):

np :)

OpenStudy (anonymous):

\[\int\limits\]

OpenStudy (yuki):

Lao, to remember the integrals of trig, you need to know the derivatives of them very well

OpenStudy (anonymous):

right ?:)

OpenStudy (yuki):

for this one, d/dx(tan(x)) = sec^2(x) was the one tat uzma referred to

OpenStudy (yuki):

@ uzma fantastic !

OpenStudy (anonymous):

thanks

OpenStudy (yuki):

another one you will see a lot is things like \[\int\limits -\csc(x)\cot(x) dx\]

OpenStudy (yuki):

if you try to do u-substitution or integration by parts, it will be a tough one, but if you remember d/dx csc(x) = -csc(x)cot(x) then we know that the integral that I showed you will have an ans csc(x) + C